OpenStudy (anonymous):

x^2-7x-8 x^2-7x-8 @Mathematics

OpenStudy (anonymous):

axis of symmetry

hero (hero):

x = -b/2a x = -(-7)/2 x = 7/2 or 3.5

OpenStudy (anonymous):

that isnt an answer choice r u sure?

hero (hero):

What are the answer choices?

hero (hero):

I'm pretty sure...

OpenStudy (anonymous):

-8,1:y=-3.5, they all have 8 in them instead of 7

hero (hero):

Strange answer choices. There's only one axis of symmetry and it's at x = 7/2 or 3.5

OpenStudy (anonymous):

identify the vertex and the axis of symmetry for the graph of y=0.4(x+3)^2-5

hero (hero):

Vertex: (-3,5) Axis Of Symmetry x = -3

OpenStudy (anonymous):

use factoring to find the zeroes of each equation? y=x^2-x-72

hero (hero):

Set y = 0, then solve for x: x^2-x-72 = 0 x^2-9x+8x-72 = 0 x(x-9)+8(x-9) = 0 (x-9)(x+8) = 0 x = -8,9

OpenStudy (anonymous):

find the zero, y=x^2+18x+81

hero (hero):

(x+9)^2 = 0 x + 9 = 0 x = -9

OpenStudy (anonymous):

find the zero, y=x^2+4x-21,and the equation of the axis of symmertry?

hero (hero):

x^2 + 7x-3x-21 = 0 x(x+7)-3(x+7) = 0 (x+7)(x-3) = 0 x = -7,3 Axis of symmetry: x = 3-(-7)/2 = 10/2 = 5

hero (hero):

-5

OpenStudy (anonymous):

which is the vertex of the graph? of y=x^2+3x-4

OpenStudy (anonymous):

wats the answer

hero (hero):

Let me verify that

OpenStudy (anonymous):

ok

hero (hero):

(-1.5, -6.25)

OpenStudy (anonymous):

find the zeroes, y=x^2+2x-4